Please help me solve the equation of 3x-9y=15.
Answer by mananth(16946)

3x-9y=15.
there is only one equation with 2 variables.
How can you solve?
At best you can put in the form of y = mx+b
..
3x-9y = 15
add 9x
3x-9y+9y=15+9y
3x=15+9y
add -15
3x-15=9y
divide by 9
3x/9 -15/9 = y
y = x/3 -5/3
...
m.ananth@hotmail.ca